OpenStudy (anonymous):

Revised: I forgot a part: f(x)=5(3)^x Find x when f(x)=105

hero (hero):

105 = 5(3)^x

hero (hero):

21 = 3^x

OpenStudy (anonymous):

lo take log on both sides to solve

hero (hero):

Yes, and x = 2.77

OpenStudy (anonymous):

Hero - How did you get the 2nd step of 21=3^x?

hero (hero):

I divided both sides by 5

OpenStudy (anonymous):

Oh!! That makes sense! Thanks!

OpenStudy (anonymous):

What about the 3rd part..how did you get your answer of 2.77?

hero (hero):

I took logs of both sides....

hero (hero):

When I did that, I got: log (21) = x log (3)

OpenStudy (anonymous):

ln(21) = ln(3^x) ln(21) = x ln(3) x = ln(21)/ln(3)

hero (hero):

Then I divided both sides by log (3): \[\frac {\log (21)} {\log (3)} = x\]

hero (hero):

2.77 = x

OpenStudy (anonymous):

Got it! Thanks!

hero (hero):

approximately